If 6 liters of a gas at a pressure of 100 kPaa are compressed reversibly according tp PV2 = C until the volume becomes 2 liters, find (a) the final pressure in kPaa, and (b) the work in kJ.
1
Expert's answer
2020-12-07T03:49:39-0500

Given V1= 6 liters

"V1= 0.006 mt^3"

P1=100 Kpa

V2= 2 liters="0.002 mt^3"

Process "P\times V^2=C"

"P1\times V1^2 = P2\times V2^2"

"100\times 6^2 = P2\times 2^2"

"P2 = 900" Kpa(final pressure)

work done = Pdv

Work done during polytrophic process n is given by

"W = \frac{P1 \times V1 - P2 \times V2}{n-1}"


"W = \frac{100\times 0.006 - 900 \times 0.002}{2-1}"


"W = - 1.2" Kj

here negative sign represents work done on the system